Sales staff for GoodsDistributors submit weekly reports listing the customer contacts made during each week. A sample of 11 weekly reports showed a sample mean of 15.3 customer contacts per week. Assume that the population of contacts is N(μ, σ²) with a sample variance of 10.89 and determine a 90% two-sided confidence interval for the population mean. Do not round any of the values when used in the final calculation. Use a DECIMAL POINT in your answer.


• z or t score = Answer


 (rounded to three decimal places)

• standard error (rounded to three decimals) = Answer



• confidence limits (rounded to one decimal) = [Answer


, Answer


] customers


1
Expert's answer
2021-11-10T07:28:56-0500

1. The critical value for "\alpha = 0.1" and "df = n-1 = 10"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.812461."


"t_c=1.812"

2.


"SE=t_c\times\dfrac{s}{\sqrt{n}}=1.812461\times\dfrac{10.89}{\sqrt{11}}\approx5.951"

standard error "=5.951"


3.


"CI=(\bar{X}-SE, \bar{X}+SE)"

"=(15.3-5.951, 15.3+5.951)"

"=(9.3, 21.3)"

confidence limits "(9.3, 21.3)" customers.